OpenStudy (anonymous):

Multiply. (2n+2)(2n-2) @Jamesshe67

OpenStudy (alexandervonhumboldt2):

differeeeence of sqrares

OpenStudy (alexandervonhumboldt2):

(a-b)*(a+b)=a^2-b^2

OpenStudy (yolo_king):

I got 4n^2-4

OpenStudy (alexandervonhumboldt2):

(2n+2)(2n-2)=4n^2-4

OpenStudy (anonymous):

(2n+2)(2n-2) =(2n+2)(2n+-2) =(2n)(2n)+(2n)(-2)+(2)(2n)+(2)(-2) =4n2-4n+4n-4 =4n2-4

OpenStudy (anonymous):

Thank you guys so much!! @YOLO_KING @AlexandervonHumboldt2 @Jamesshe67 you guys down to help with some more ?:)

OpenStudy (anonymous):

yup

OpenStudy (yolo_king):

Sure C:

OpenStudy (anonymous):

Cool Cool!

OpenStudy (alexandervonhumboldt2):

yes can you medal me andd fan us

OpenStudy (yolo_king):

C:^^^^

OpenStudy (anonymous):

Yes but you guys all gave me answers so i have to do it fairly ;) Help me on the next ones and you will all get medals . promise!